OpenStudy (anonymous):

A tennis ball connected to a string is spun around in a vertical, circular path at a uniform speed. The ball has a mass m = 0.16 kg and moves at v = 5.33 m/s. The circular path has a radius of R = 1.13 m What is the minimum velocity so the string will not go slack as the ball moves around the circle?

OpenStudy (anonymous):

You need to find the centripetal acceleration that is at least equal to the acceleration of gravity.

OpenStudy (anonymous):



OpenStudy (anonymous):

once i find the centripital accel then what?

OpenStudy (anonymous):

mg = mV^2/r g=V^2/r sqrt(rg) =V

OpenStudy (anonymous):

thank you soo much!

OpenStudy (perl):

@CliffSedge Your tension vector is not pointing in the right direction, tension is pointing towards the center

OpenStudy (anonymous):

@perl. Incorrect. Tension is pointing towards the center from the ball, but tension is also pointing towards the ball from the center (Newton's Third Law). It is just a schematic of the situation and not a free-body diagram.